If M/N+N/L+L/M=0 then find out LN/M square+M square/LN -N cube/L cube= ?

Grade:12th pass

1 Answers

Ram Kushwah
3 years ago
GIVEN THAT
 
M / N + N / L + L / M =0
 
LM² + MN²  + L²N=0
 
LM²  + L²N= - MN²
 
SUARING BOTH SIDES
 
(LM²  + L²N )² = (- MN² )²
 
L²M4 + L4N²  +2 L³M²N = M²N4
 
L²M4 + L4N²- M²N4  = – 2 L³M²N
 
DIVIDING BOTH SIDED BY  L³M²N
 
L²M4 /  L³M²N + L4N² / L³M²N  –  M²N4 /  L³M²N  = – 2 L³M²N / L³M²N
 
M²/ LN  + LN / M²  – N³ / L³ = -2
 
OR  LN / M² + M² / LN – N³ / L³ = – 2
 
HENCE THE FINAL ANSWER = – 2